Please provide detailed step by step solution to this problem. Thank you in advance !!

The blades of a fan running at low speed turn at 210 rpm. When the fan is switched to high speed, the rotation rate increases uniformly to 440 rpm in 5.10 s. What is the magnitude of the angular acceleration of the blades ? Answer in rad/s²
Answer #1

The angular acceleration = (440 - 210)rpm/5.1 s
= 230/ (60*5.1) rev/s^2
= 0.752 rev/s^2 = 4.72 rad/s^2